### v2

# EmMixformer: Mix transformer for eye movement recognition

Abstract
Eye movement (EM) is a new highly secure biometric behavioral modality that has received increasing attention in recent years. Although deep neural networks, such as convolutional neural network (CNN), have recently achieved promising performance, current solutions fail to capture local and global temporal dependencies within eye movement data. To overcome this problem, we propose in this paper a mixed transformer termed EmMixformer to extract time and frequency domain information for eye movement recognition. To this end, we propose a mixed block consisting of three modules, transformer, attention Long short-term memory (attention LSTM), and Fourier transformer. We are the first to attempt leveraging transformer to learn long temporal dependencies within eye movement. Second, we incorporate the attention mechanism into LSTM to propose attention LSTM with the aim to learn short temporal dependencies. Third, we perform self attention in the frequency domain to learn global features. As the three modules provide complementary feature representations in terms of local and global dependencies, the proposed EmMixformer is capable of improving recognition accuracy. The experimental results on our eye movement dataset and two public eye movement datasets show that the proposed EmMixformer outperforms the state of the art by achieving the lowest verification error.
Index Terms:
Biometrics, Eye movements, Transformer, LSTM, Attention
I Introduction
Biometrics refers to identifying or verifying person’s identity based on their biometric traits[1]. Unlike traditional non-biometric methods such as key, password, and card, biometric traits are difficult to copy and forge, and cannot be forgotten, making them highly secure and convenient. Current biometric authentication techniques can be categorized into two types: Physiological and Behavioral modalities [2]. The former utilize static body biometrics, such as fingerprint recognition [3], face recognition [4], vein recognition [5], and iris recognition[6], while the latter are related to user dynamic behavior, such as eye movements[7], gait [8], signature[9], and voice[10]. Physiological modalities are intrusive and may require user’s cooperation, which may be inconvenient. Additionally, they are prone to being recreated or captured, in addition to being vulnerable to spoofing as they can be artificially synthesized by using, for instance, gummy fingers[11] or deepfake technology. Dynamic biometric traits, by contrast, exploring active and dynamic personal characteristics, are unique and almost impossible to spoof since they are caused by subconscious processes. Recently, behavioral biometrics, such as eye movement with temporal information associated with human cognition, have received increasing focus in biometrics.
I-A The principle of Eye movement
Eye movement biometrics [7] aims to identify the behavioral patterns and information regarding physiological properties of tissues and muscles generating eye movement. These behavioral and physiological patterns provide rich information about the cognitive brain functions and neural signals controlling eye movements. Compared to other biometrics, eye movement biometrics technology has significant advantages. It not only provides liveness detection [12], ensuring thereby security even in unconscious states, but its reliance on both the properties of the oculomotor plant and the neurological properties controlled by the brainstem makes eye movements almost impossible to replicate [13]. Eye movement biometrics, in addition, are highly suitable for multi-modality recognition as it may be seamlessly combined with other eye-related physiological traits such as iris and pupil[14]. These key advantages explain why eye movement biometrics has become a hot research topic over the recent years.
The human eye moves within six degrees of freedom. Oculomotor movements encompass the coordinated contractions and relaxations of the six eye muscles (four rectus and two obliques), controlled by the brainstem nerves during visual processing (Fig.1). These movements serve as a vital mechanism for humans to selectively extract information from the visual environment. As they enable us to focus our attention on objects of interest, eye movements, like keystroke dynamics[15], gait [8] and electrocardiography (ECG)[16], contain rich information about cognitive functions, neural pathways, and other brain-related aspects [17]. The literature shows that the way we move our eyes in response to a given stimulus is highly individual [18, 19] and that these individual characteristics are reliable over time [20]. Hence, eye movements have been proposed as a behavioral biometrics [21, 22].
Eye movements can be divided into saccades and fixations [23]. Saccades involve rapid eye movements as the gaze transitions from one point to another, typically occurring at speeds of 200-400°/s, and peak velocities that may exceed 600°/s[24]. To initiate a saccade, related neural circuitry calculates the disparity between the starting and target positions. It then sends precise neural signals to the extra-ocular muscles, which in turn rotate the eye accordingly. If the eyeball fails to accurately reach the intended target, the muscles will make minute corrections. Fixations, by contrast, relatively focus on a specific gaze point while exhibiting high-frequency, low-amplitude nystagmus movements, and low-velocity eye movements with a typical duration of 100-400 ms. Eye tracking devices can capture both position and movement data for saccades and fixations, which can be leveraged to extract eye movement features for biometric identification.
Figure 1: Eye Musculature.
I-B Related Work
Traditional eye movement recognition approaches
Traditional approaches use handcrafted descriptors and machine learning (ML) algorithms for feature extraction and recognition. Kasprowski and Ober[21] made a first attempt to explore eye movement features for identification. Their experimental results on a small dataset showed that eye movement comprises discriminating information, allowing the classification model to achieve low verification error. In 2011, Holland and Komogortsev [25] gathered eye-movement scan path data generated by 32 volunteers and designed 15 eye movement metrics for feature extraction. A weighted average method was then used to improve accuracy. In 2012, Komogortsev et al. [26] introduced the concept of oculomotor plant characteristics (OPC) for identification purposes and employed Student’s t-test enhanced by voting and Hotelling’s T-Square test to distinguish between gaze and sweep events. Subsequently, they [13] extracted more intricate features such as eye hopping abnormalities, compound eye hopping, and dynamic overshooting from eye movements, input to Random Forest (RF) and Support Vector Machines (SVM) for classification. In 2017, Bayat and Pomplun [27] extracted a feature set comprising 28 features for identification by using gaze, sweep, and pupil size information. They combined three classifiers and evaluated performance on a dataset of 40 objects. Similarly, Cuong et al. [28] proposed a method to encode eye movement using Mel-Frequency Cepstrum Coefficients (MFCCs) without distinguishing between sweeping and gazing. In 2018, Li et al. [29] proposed a multi-channel Gabor wavelet transform to extract texture features from eye movement trajectories and SVM was employed for recognition. In [30], Lohr et al. proposed three multilayer perceptrons (MLP) with triplet loss to learn the features of three event types, fixation, saccade, and post-saccadic oscillations (PSO). In recent years, Long Short-Term Memory (LSTM) was proposed to extract temporal features for gaze pattern recognition[31, 32].
Deep Learning based eye movement recognition approaches
Recently, deep learning (DL) models, e.g. Convolutional Neural Networks (CNNs), thanks to their robust feature representation, were successfully applied for image segmentation[33], image classification[34], object detection[35], and, recently, eye movement recognition. Lohr et al. [7], for instance, introduced an exponentially-dilated CNN with multi-similarity loss for user authentication. Taha et al. [36] combined LSTM and dense networks to learn temporal characteristics from the eye movement (EM) profiles and proposed an end-to-end learning model for driver authentication. Likewise, [37] developed a CNN to learn feature representation from the raw eye-tracking signal. For the same purpose, Makowski et al. [38] investigated a CNN to process binocular eye tracking signals and the relative positions of the stimuli to detect replay attacks. In 2022, Makowski et al. [39] collected an eye tracking dataset from 66 participants in sober, fatigued and alcohol-intoxicated states, and fine-tuned the DeepEyedentification [37] model to investigate the effects of drunkenness and fatigue on eye movement biometrics. At the same year, D. Lohr et al. [40] proposed a DenseNet network architecture for end-to-end eye movement biometric authentication.
I-C Motivation and Contribution
Traditional approaches use manually-designed algorithms to extract, from eye movement, handcrafted features, input to ML approaches such as SVM and MLP, which suffers from the following issues: 1) Manually-designed models rely on human prior assumptions such as Gaussian distributions. These assumptions, however, are not always effective to learn robust eye movement patterns as the latter’s underlying data distributions are usually much more complex in real-life applications. This explains also why it is hard to design a mathematical model to effectively model such distributions; 2) Traditional ML approaches such as SVM and RF can be viewed as a neural network with very few (one or two) layers for feature representation learning. Their representation capacity is therefore relatively limited. DL models trained on huge data, by contrast, are capable of learning high-level features by stacking a number of layers. For eye movement recognition, DL models, such as CNN, have recently achieved promising performance compared to ML approaches and become an increasing modeling trend. CNN can effectively capture local patterns and features, such as edges, corners, textures, or shapes, by sharing weights and local perception. CNN, however, is difficult to capture long temporal dependencies within sequences because of limited receptive field[41][42]. As eye movement is typically temporal data, collected over time with observations recorded at different time points, existing DL-based approaches is difficult to extract effective features that embed such long-range time dependencies. Recently, transformers with attention mechanisms have emerged as new start-of-the art classifiers thanks to their high capacity to capture long dependencies among tokens, with impressive performance obtained, first, in various natural language processing (NLP) tasks [42] and then in computer vision [43]. In fact, the capacity of model to learn long dependencies is related to the length of the paths forward and backward signals. The shorter paths between any combination of positions in the input and output sequences allow to learn longer range dependencies[44]. For example, the work [42] compares the maximum path length between any two input and output positions in networks such as transformer, LSTM, and CNN and shows that transformer has the shortest paths and is capable of learning long range dependency.
Inspired by these facts, we propose in this paper a novel mix transformer for eye movement recognition, named EmMixformer. EmMixformer includes three modules, transformer, attention-LSTM, and Fourier-transformer. The eye movement signal is first split into fast and slow velocity signals, separately input to encoder for feature embedding. The embedding features are then input to the mix transformer module to learn a robust representation for classification. The transformer and attention LSTM aims at learning temporal dependencies within the eye movement signal, while the Fourier transformer learns a global feature representation. Our mix transformer can, therefore, learn complementary features. The contributions of our work are summarized as follows:
  • We are the first to propose transformer for the eye movement recognition task by proposing an eye movement mix transformer (EmMixformer) to learn a robust feature representation. Concretely, we investigate a mix block, consisting of three modules (as shown in Fig.2): a transformer, an attention LSTM and a Fourier-transformer. Our approach is capable of extracting eye movement features in both time and frequency domains.
  • We design an attention LSTM to learn robust temporal dependencies by incorporating the attention mechanism into LSTM structure. Our resulting attention-based LSTM model is capable of capturing robust features embedding temporal dependencies for eye movement recognition.
  • We propose a Fourier transformer to learn global dependencies by performing self attention in the frequency domain. This is the first transformer to perform frequency domain learning.
  • We build a low-frequency eye tracking dataset acquired thanks to a state-of-the-art portable eye-tracking instrument (Tobii Pro glasses3) and make it publicly available to promote research on eye movement recognition. The data captured by this low cost sensor are more realistic, which promotes more valuable benchmarking in the area of eye movement recognition.
  • We have carried out rigorous experiments on our eye movement dataset and on two public eye movement datasets to assess our approach. The experimental results show that our approach outperforms existing works in terms of reducing the verification error, and achieves new state-of-the-art-recognition results.
II Methodology
In this section, we detail the proposed EmMixformer model, as illustrated in Fig.2. EmMixformer comprises a prepossessing block, a Siamese CNN block, and a Mix block. The prepossessing blorck splits the input eye movement data into low and fast velocity data. The Siamese CNN block comprises two sub-nets for feature embedding of the two velocity data, while the Mix block comprises three modules, transformer, attention LSTM (attLSTM) and Fourier transformer (Fourierformer). The slow and fast velocity signals output by the prepossessing block are input to the Siamese CNN block for feature embedding. The resulting features are forwarded to the mix block for feature extraction and recognition. We detail, in this section, the Siamese CNN block and the Mix block, and provide the details of the prepossessing block in Section III.
Figure 2: The framework of the proposed EmMixformer model
II-A Siamese CNN block
The Siamese CNN block consists of two sub-networks, each stacking four one-dimensional convolutional modules for feature extraction. The size of the convolution kernel increases to obtain a progressively larger receptive field. Each module comprises a convolutional layer, a batch normalization layer, a ReLU layer, and an average pooling layer. The two subnets take respectively the preprocessed fast and slow eye movement signals as inputs and output the corresponding embedding feature representations. We then concatenate the output feature representations from the two subnets to obtain the complete feature representation. The $x$ and $y$ represent the horizontal and vertical coordinate sequences of eye movement data, and $n$ represents the sequence length.
II-B Mix block
The Mix block comprises three modules, an attention LSTM (attLSTM), a Transformer, and a Fourier-transformer (Fourierformer). The three modules take the Siamese CNN block’s output as input to extract rich and comprehensive temporal-spatial features. Concretely, a temporal sequence from the Siamese CNN block is divided into a number of sub-sequences, which are input to attLSTM and transformer to extract the temporal features. The attLSTM module learns short-term dependencies within the temporal sequences, while the transformer module extracts long-term dependencies within the temporal sequences thanks to its self-attention mechanism [45]. Their combination, therefore, can learn a comprehensive temporal feature representation. The third module, Fourierformer, takes as input the temporal sequences to learn a global frequency-domain feature representation. As a result, our mix block extracts both a feature representation embedding temporal dependencies and a global feature representation, thus effectively improving the modeling of time-series eye movement classification.
Transformer
Transformer, introduced in 2017 by Vaswani et al [42], has become the state of the art for NLP [42] and more recently computer vision [45]. Unlike traditional sequence models such as recurrent neural networks (RNNs) or CNNs, Transformer relies solely on a key idea, the self-attention mechanism, eliminating the need for sequential processing. The self-attention mechanism allows the model to weigh the importance of different tokens in a sequence, and enables it to capture effectively long-range dependencies and context information. By attending to all tokens simultaneously, Transformer can process sentences in parallel, leading to significantly faster training and inference processing times, compared to existing sequential models [45].
For eye movement sequence $X$=($x_{1}$,…, $x_{T}$), position encoding is introduced by adding the position information into the embedding vector. We treat each element in the eye movement sequence as a token, which yields $T$ tokens for $X$. To learn the temporal feature representation, we perform self-attention among tokens. We transform the $t$th token $x_{t}$ into three vectors, query $q_{t}$, key $k_{t}$, and value $v_{t}$, by three weight matrices $w_{q}$, $w_{k}$, and $w_{v}$. We pack the $T$ query vectors $q_{1},q_{2},...,q_{T}$ of the input sequence into a matrix $Q$. Similarly, the keys and values are packed together to obtain matrix $K$ and $V$. The self-attention among $T$ tokens is implemented by Eq.(1).
$Attention\left(Q,K,V\right)=Softmax\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
where Softmax($\cdot$) is applied to the rows of similarity matrix $QK^{T}$, and $d_{k}$ provides normalization. Query $Q$, keys $K$, and values $V$ are the projections of the tokens computed by projection matrices $W_{q}$, $W_{k}$, and $W_{v}$, respectively.
To learn a robust feature representation, we set a multi-head self attention by concatenating $L$ attention heads (Eq.(2)).
$MultiHead=Concat\left(head_{1},...,head_{L}\right)W$ (2)
where $L=4$, $head_{l}=Softmax\left(\frac{Q_{l}K_{l}^{T}}{\sqrt{d^{\prime}_{k}}}\right)V_{l}$ denotes the output of the $l$th attention head, and $W$ is the linear transformation matrix. Subsequently, the representation $MultiHead$ is subjected to residual concatenation and layer normalization operations by Eq.(3):
$Z=\left(X+MultiHead(LN(X))\right)$ (3)
Finally the representation of each node is non-linearly transformed by a feed-forward neural network with two layers:
$X_{transformer}=LN(MLP(LN(Z))+Z)$ (4)
To simplify representation, we combine Eq.(1), Eq.(2), Eq.(3), Eq.( 4) to obtain Eq.( 5).
$X_{transformer}=transformer(X)$ (5)
Attention LSTM
[…]
$f_{t}=\sigma\left(W_{f}{x}+U_{f}h_{t-1}+V_{f}C_{t-1}+b_{f}\right)$ (6)
$i_{t}=\sigma\left(W_{i}{x}+U_{i}h_{t-1}+V_{i}C_{t-1}+b_{i}\right)$ (7)
$o_{t}=\sigma\left(W_{o}{x}+U_{o}h_{t-1}+V_{o}C_{t}+b_{o}\right)$ (8)
$\widetilde{C_{t}}=tanh\left(W_{C}{x}+U_{C}h_{t-1}+b_{C}\right)$ (9)
$C_{t}=f_{t}\odot C_{t-1}+i_{t}\odot\widetilde{C_{t}}$ (10)
$h_{t}=o_{t}*tanh({C_{t}})$ (11)
where $f_{t}$ , $i_{t}$ and $o_{t}$ are the forgetting gate, input gate and output gate, respectively. $\widetilde{C_{t}}$ , $C_{t}$ and $h_{t-1}$ are the candidate cell state, the cell state, and the hidden state at the last moment, respectively. $\odot$ denotes the element-by-element matrix multiplication and $\sigma$ represents the sigmoid activation function, respectively.
Figure 3: Attention LSTM
To improve LSTM learning capacity, we incorporate the attention mechanism to propose attention LSTM, as shown in Fig. 3. Let $x$, $h^{\prime}_{t-1}$ and $C^{\prime}_{t-1}$ be the input vector, previous hidden state and previous cell state in our attention LSTM. We calculate query matrix $Q_{x}$, key matrix $K_{x}$, and value matrix $V_{x}$ of $x_{t}$. Similarly, we obtain query matrix $Q_{h}$, key matrix $K_{h}$, and value matrix $V_{h}$ of $h_{t-1}$, and query matrix $Q_{c}$, key matrix $K_{c}$, and value matrix $V_{c}$ of $C_{t-1}$. Based on Eq(1), the self attention of input vector $x_{t}$ is computed by Eq. (12):
$\begin{gathered}SA_{x}=Attention(Q_{x},K_{x},V_{x})\\ \end{gathered}$ (12)
Similarly, the self attention of hidden state $h_{t-1}$, and the cross attention between $x_{t}$ and $h_{t-1}$ are calculated by:
$\begin{gathered}SA_{h}=Attention(Q_{h},K_{h},V_{h}),\end{gathered}$ (13)
$\begin{gathered}CA_{x}=Attention(Q_{x},K_{h},V_{h}),\end{gathered}$ (14)
$\begin{gathered}CA_{h}=Attention(Q_{h},K_{x},V_{x}),\end{gathered}$ (15)
$CA_{c}=Attention(Q_{c},K_{h},V_{h}),$ (16)
where $CA_{x}$ and $CA_{h}$ denote cross attention between $x_{t}$ and $h_{t-1}$, $CA_{c}$ is the cross-attention vector of cell state $C_{t-1}$ and hidden state $h_{t-1}$, while $SA_{h}$ and $SA_{x}$ represents the self attention of the hidden state and input.
The four gates in our attention LSTM are computed by:
$i^{\prime}_{t}=\sigma\left(linear(Concat\left(SA_{x},SA_{h},CA_{c}\right))\right)$ (17)
$f^{\prime}_{t}=\sigma\left(linear(Concat\left(CA_{h},SA_{x},CA_{c}\right))\right)$ (18)
$o^{\prime}_{t}=\sigma\left(linear(Concat\left(CA_{x},SA_{h}\right))\right)$ (19)
$\widetilde{C^{\prime}_{t}}=tanh\left(linear(Concat\left(SA_{x},SA_{h}\right))\right)$ (20)
$C^{\prime}_{t}=f^{\prime}_{t}\odot C^{\prime}_{t-1}+i_{t}\odot\widetilde{C^{\prime}_{t}}$ (21)
$h^{\prime}_{t}=o^{\prime}_{t}*tanh({C^{\prime}_{t}})$ (22)
The LSTM takes sequence ($x_{1}$,…, $x_{t}$) with 256 dimensions and outputs hidden state sequence ($h^{\prime}_{1}$,…, $h^{\prime}_{T}$). To simplify the description, we pack the Eq.(12-21) into Eq.(23):
$X_{attLSTM}=attLSTM\left(x_{1},x_{2},...,x_{t}\right)$ (23)
Fourier transformer
To learn a global feature representation, we perform self attention in the frequency domain by Fourier transformation (Fig.4). Based on Fourier theory [51] [52], feature learning in frequency domain has an image-wide receptive field by channel-wise Fourier transformation.
Figure 4: Fourier transformer
A one-dimensional discrete Fourier transform (DFT) transforms sequence $X$ from time domain to frequency domain (Eq.(24)):
$\mathcal{F}(X)\left(k\right)=\frac{1}{\sqrt{T}}\sum_{t=0}^{T-1}X(t)e^{-j\frac{2\pi}{T}kt}$ (24)
where $\mathcal{F}(X)\left(k\right)$ is the complex value of the $k$-th frequency component in the frequency domain, $X(t)$ is the complex value of the $t$th sampling point in the time domain, $T$ is the sequence length, and $\frac{1}{\sqrt{T}}$ ensures normalized consistency between the DFT and the inverse Fourier transform (IDFT).
Using the real part $Re(X)$ and the imaginary part $Im(X)$ of $\mathcal{F}(X)\left(k\right)$, we obtain the amplitude component $A(X)$ and the phase component $\varphi\left(X\right)$ by the following equations:
$A(X)\left(v\right)=\sqrt{Re\left(X\right)\left(k\right)^{2}+Im\left(X\right)\left(k\right)^{2}}$ (25)
$\varphi\left(X\right)\left(k\right)=arctan\left(\frac{Im\left(X\right)\left(k\right)}{Re\left(X\right)\left(k\right)}\right)$ (26)
Next, we extract the feature representation from amplitude component $A(X)\left(k\right)$ and phase component $\varphi(X)\left(k\right)$ using two independent transformers. The self attention operation is performed to capture the amplitude information $A_{transformer}$ and phase information $\varphi_{transformer}$, computed as:
$A_{transformer}=transformer(A(X))$ (27)
$\varphi_{transformer}=transformer(\varphi(X))$ (28)
Subsequently, the feature representations of the amplitude and phase components are combined into complex numbers and transformed back into the spatial domain by IFFT
$X^{\prime}\left(t\right)=\mathcal{F}^{-1}\left(complex(A_{transformer},\varphi_{transformer})\right)$ (29)
Finally, we concatenate the outputs of the three modules, i.e. attention LSTM, Transformer, and Fourier transformer, to obtain the mix block (Eq.(30)), with the mixing block stacked in two layers.
$Y=Concat(X_{attLSTM},X_{transformer},X^{\prime})$ (30)
III Experimental results
To assess our approach, we have carried out extensive experiments on our self-built dataset and two public datasets, GazeBase [53] and JuDo1000 [38]. We compare our approach with state-of-the-art works, namely DEL[38], Expansion CNN[7], Dense LSTM[36] and DenseNet [40]. All experiments were conducted in Pytorch and implemented on a high performance computer with a NVIDIA GTX 3090Ti GPU.
III-A Dataset
TABLE I: Various eye movement datasets
Data set Device Freq.(Hz) Chin rest Subj.
EMglasses Tobii pro glasses3 50 no 203
JuDo1000[38] EyeLink Portable 1000 yes 150
GazeBase[53] EyeLink 1000 1000 yes 322
CEM-{CJK*}UTF8gbsnⅠ[23] Tobii TX300 300 yes 22
CEM-{CJK*}UTF8gbsnⅡ[23] EyeLink 1000 1000 yes 32
VREM[54] ET-HMD 250 yes 68
SBA-ST[55] EyeLink 1000 1000 yes 298
EMglasses
[…]
Figure 5: Data collecting
In each experiment, a black-white ball with a diameter of 1 cm, displayed as a stimulus, was randomly moved at the screen and stopped for 1 second in each position. Participants tracked the ball for about 10 seconds by moving their gaze point. During each recording, the gaze location was monitored by the proctor to ensure compliance. As shown in Fig.6, the red circle represents the position of the volunteer’s gaze point, while the curves below are the time-series visualization displays of eye movement. Participants completed data collection from two sessions, with the proctor suggesting to participants to take a few minutes break between each session if needed. As a result, 406 (2 sessions $\times$ 203 participants) sequences from 203 participants were collected.
Figure 6: Collected data.
GazeBase
The GazeBase dataset [53] comprises eye-movement recordings from 322 volunteers using the EyeLink1000 eye-tracker instrument. Volunteers’ heads and chins were immobilized at a distance of 550 mm from a 47.4 $\times$ 29.7 cm screen during data acquisition. Each volunteer was exposed to seven tasks with different stimulus materials during acquisition, namely a randomized scanning task (RAN), a reading task (TEX), two video viewing tasks (VD1 and VD2), a gaze task (FXS), a horizontal scanning task (HSS), and an eye-driven game task (BLG). Nine recording rounds were conducted over 37 months, where the participants in each round were recruited exclusively from the previous round. In our experiments, four sub-datasets in GazeBase, RAN, TEX, FXS, and HSS, are used for recognition.
JuDo1000
The JuDo1000 dataset [38] contains eye movement data from 150 volunteers, each participating in four experimental sessions with a time interval of more than one week between any two sessions. The data were recorded using an EyeLink Portable Duo eye tracker with a sampling frequency of 1,000 Hz. The stimulus was a black dot with a diameter of 0.59 cm that appeared at five random locations on the screen during the experiment, with duration varying between 250, 500, and 1000 ms. Participants were seated in front of a 38×30 cm computer monitor at a height adjustable table and a viewing distance of 68 cm, with their head stabilized by a chin and forehead rest stand.
III-B Signal preprocessing
As the scope of saccadic and fixational velocity are vastly different, global normalization would make the slow fixational drift and tremor to near zero, which results in missing key information in the eye tracking signal. Similar to works [38][7], we split the original data into two subnets: a fast subnet and a slow subnet. Let X = ($x_{1},...,x_{t}$) and Y = ($y_{1},...,y_{t}$) be eye gaze point coordinate data. First, we transform the coordinate information into velocity information at each sampling point using following equations :
$\delta^{(i)}_{x}=\frac{x^{(i)}-x^{(i-1)}}{t^{(i)}-t^{(i-1)}}$ (31)
$\delta^{(i)}_{y}=\frac{y^{(i)}-y^{(i-1)}}{t^{(i)}-t^{(i-1)}}$ (32)
Then, we remove points without gaze data and replace the Nan values with 0. As a result, each eye data includes two channels: $\delta_{x}$ for the horizontal velocity sequence and $\delta_{y}$ for the vertical velocity sequence.
Furthermore, the preprocessing approach [38] [7] is used to convert the raw data into two parts: fast data $D_{fast}$ and slow data $D_{slow}$. For the fast data, absolute velocities below a minimal velocity $v_{min}$ are truncated and z-score normalization is employed for normalization.
$D_{fast}^{(i)}{(\delta^{i}_{x},\delta^{i}_{y})}=\left\{\begin{matrix}Z(\delta^{i}_{x}),Z(\delta^{i}_{y})\quad\quad\quad otherwise\\ \quad Z(0)\quad\quad if\sqrt{(\delta^{i}_{x})^{2}+(\delta^{i}_{y})^{2}}<v_{min}\end{matrix}\right.$ (33)
$Z_{i}=(\delta^{i}-\mu)/\sigma$ (34)
where $\delta^{i}$ is the velocity at the $i$-th sampling point, and $\mu$ and $\sigma$ are the mean and variance of the sequence $\delta$.
For slow data $D_{slow}$, we use the hyperbolic tangent function to transform the input into a range of (-1, +1):
$D_{slow}^{(i)}{(\delta^{i}_{x},\delta^{i}_{y})}=(tanh(c\delta^{i}_{x}),tanh(c\delta^{i}_{y}))$ (35)
The velocity threshold $v_{min}$ = 40 and the scaling factor c = 0.02 are two fixed hyperparameters.
After preprocessing, the fast data and slow data are input to Siamese CNN for feature embedding (as shown in Fig.2). The latter’s outputs are further input to the mix block for feature extraction and classification.
III-C Experimental settings
To assess our approach, we split each dataset into training and testing sets. In our dataset, comprising 406 sequences from 203 participants, we select data collected at the first session for training and the remaining data for testing. For the GazeBase dataset, we consider the four sub-datasets, RAN, TEX, FXS, and HSS, with each providing 9 rounds (Rounds 1-9) of data collected over 37-months, and two collections per round. As participants in a round were recruited only from the previous round, the first round collection includes data from all participants. After Round 6, fewer than 60 participants remain. The data from the first and second sessions in the first round are considered for training and testing, respectively. As a result, the training and testing sets both include 322 sequences for the first round. Accordingly, the four sub-datasets, RAN, TEX, FXS, and HSS, result in four training sets and testing sets at the first round. To evaluate the generalization ability of our approach w.r.t processed data at long time intervals, round 1-6 data from the RAN dataset (as listed in Table II) is also considered for comparable experiments. Specifically, we used round 1 data collected from the first session as the training set and round 1-6 data collected from the second session as the testing set. The training and test sets both include 322 sequences (322 participants $\times$ 2) at the first round. At the second round, we have 136 participants that had also provided the eye movement data at the first round. So, we select the data from these 136 participants at the first round for training and the data from the same 136 participants at the second round for testing. Accordingly, the training and the testing sets both include 136 sequences. In the same way, we obtain the training and testing sets for the remaining 3-6 rounds. Table II shows information regarding the RAN sub-dataset.
TABLE II: The RAN database collects information from the first round to the sixth round.
Information Round1 Round2 Round3 Round4 Round5 Round6
Time interval(month) 0 1 2 2 9 15
Number of participants 322 136 105 101 78 59
For JuDo1000, 600 sequences were collected from 150 participants, each providing data for four times. The data from the first three times are selected for training while the last collection is selected as a test set. Accordingly, there are 450 sequences (150 participants $\times$ 3 times) in the training set and 150 sequences (150 participants $\times$ 1 times) in the testing set.
In our dataset, 203 participants provided eye movement data from two sessions with few minutes interval at one round. The data from the first session are used for training and the data from the second session for testing. Accordingly, the training and testing sets both comprise 203 sequences.
To optimize the network parameters, we used the cross-entropy loss function in conjunction with the Adam optimizer. The cross-entropy loss was used to quantify the disparity between the model’s output and the actual labels, enabling thereby the model to learn correct classifications. The Adam optimizer dynamically adjusts the network parameters to continuously enhance model’s performance during training. The learning rate was fixed to 0.0002 and the batch size set to 64. All models were trained for 1000 rounds.
To assess our approach, the Equal Error Rate (EER) is computed. EER indicates the error rate when the False Acceptance Rate (FAR) and False Rejection Rate (FRR) are equal. Lower EER values indicate better verification performance. We also report the FRR at different FARs $10^{-1}$, $10^{-2}$, and $10^{-3}$.
III-D Experimental results for short time interval
These experiments aim to assess the verification performance of our approach on short time interval data. In the three datasets, the participants provided data from two sessions at each round. As detailed in section III-C, we select data from the first session for training and the remaining data for testing. The EMglasses dataset comprises 203 sequences in the training set and testing set, the four sub-datasets (RAN, TEX, FXS, and HSS) in the the GazeBase dataset comprises 322 sequences in the training set and the testing set, while the JuDo1000 dataset comprises 450 sequences in the training set and 150 sequences in the testing set. Note that in all datasets, we split each sequence, collected for several seconds, into several sub-sequences for training and testing. The state of the art approaches, i.e. DEL[38], Expansion CNN[7], Dense LSTM[36] and DenseNet [40] are employed for benchmarking. The EER and FRR@FAR of each approach are listed in Table III, Table IV, Table VII, Table V, Table VI, and Table VIII. The corresponding ROC curves are shown in Fig. 7. It can be observed that our EmMixFormer outperforms the existing approaches and achieves the lowest verification error, i.e. 0.0673, 0.0635, 0.0801, 0.1578 on the four subsets of the GazeBase dataset, 0.0502 on the JuDo1000 dataset, and 0.1599 on our dataset. Also, we observe from Fig.7 that our proposed approach achieves the higher recognition accuracy w.r.t existing approaches at different FARs.
TABLE III: Results of comparative experiments on the HSS database
HSS EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
EmMixformer 0.0673 0.0502 0.2032 0.4659
DEL[38] 0.1309 0.1680 0.6217 0.9087
Expansion CNN[7] 0.1437 0.1894 0.6267 0.9245
Dense LSTM[36] 0.1191 0.1365 0.5082 0.8382
DenseNet[40] 0.0839 0.0739 0.2691 0.5575
TABLE IV: Results of comparative experiments on the TEX database
TEX EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
EmMixformer 0.0635 0.0407 0.2603 0.6193
DEL[38] 0.1060 0.1128 0.5750 0.9141
Expansion CNN[7] 0.1362 0.1950 0.6977 1.0000
Dense LSTM[36] 0.0971 0.0945 0.4824 0.8456
DenseNet[40] 0.0736 0.0551 0.3293 0.7175
TABLE V: Results of comparative experiments on the RAN database
RAN EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
EmMixformer 0.0801 0.0680 0.2818 0.6087
DEL[38] 0.1436 0.2066 0.7383 0.9645
Expansion CNN[7] 0.1500 0.2340 0.7277 0.9902
Dense LSTM[36] 0.1161 0.1329 0.5529 0.8846
DenseNet[40] 0.0885 0.0807 0.3513 0.7045
TABLE VI: Results of comparative experiments on the FXS database
FXS EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
EmMixformer 0.1578 0.2549 0.7654 0.9539
DEL[38] 0.2229 0.5022 0.9358 0.9936
Expansion CNN[7] 0.1950 0.4051 0.9067 0.9903
Dense LSTM[36] 0.2191 0.4047 0.8831 0.9832
DenseNet[40] 0.1666 0.2979 0.8323 0.9821
TABLE VII: Results of comparative experiments on the JuDo1000 database
JuDo1000 EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
EmMixformer 0.0543 0.0059 0.1284 0.3359
DEL[38] 0.1238 0.0781 0.5508 0.8945
Expansion CNN[7] 0.0989 0.0586 0.3203 0.7594
Dense LSTM[36] 0.0669 0.0195 0.2305 0.6016
DenseNet[40] 0.0773 0.0125 0.1953 0.4922
TABLE VIII: Results of comparative experiments on the EMglasses database
EMglasses EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
EmMixformer 0.1599 0.2572 0.7888 0.9711
DEL[38] 0.1853 0.3565 0.8877 0.9900
Expansion CNN[7] 0.2140 0.4013 0.8654 0.9821
Dense LSTM[36] 0.2069 0.4545 0.8928 1.0000
DenseNet[40] 0.1892 0.3444 0.8131 0.9812
Figure 7: ROC Curve of various approaches on (a) HSS, (b) TEX, (c) RAN, (d) FXS in the GazeBase dataset, (e) JuDo1000 dataset, and (d) our dataset
III-E Experimental results for long time intervals
In the comparative experiments, we evaluated the authentication performance of our model on eye-tracking data collected at long time intervals. The RAN sub-dataset in GazeBase comprises 1-6 round eye movement data over a 37-month collection cycle with the users providing data from two sessions for each round. As detailed in Section III-C, data collected at the first session in round 1 are used for model training, while data collected at the second session from round 1 to round 6 are used for test. Detailed information of Round 1-6 in the RAN dataset is given in Table II, where the time interval represents the duration in months of each round. The verification error rates are listed in Table IX. We observe that our EmMixFormer consistently achieves the lowest EER for the different datasets collected at different times. It is worth noting also that all methods obtain higher verification error rates with increasing time intervals.
TABLE IX: Results of long-term interval experiments on the RAN database
Model EER
Round1 Round2 Round3 Round4 Round5 Round6
EmMixformer 0.0801 0.1428 0.1803 0.1793 0.1853 0.2279
DEL[38] 0.1436 0.2215 0.2594 0.2926 0.2459 0.2907
Expansion CNN[7] 0.1500 0.2328 0.2769 0.2395 0.2812 0.3057
Dense LSTM[36] 0.1161 0.1939 0.2535 0.2791 0.2322 0.2840
DenseNet[40] 0.0885 0.1775 0.2161 0.2032 0.2047 0.2592
Overall, all the experimental results show that the proposed EmMixFormer significantly outperforms existing eye movement recognition approaches and achieves the lowest verification errors on the three datasets. Such good performance is explained by the following facts. The CNN based approaches, i.e. DEL[38], Expansion CNN[7], and DenseNet[40], leverage the local receptive fields and shared weights to learn local structure, but are unable to learn robust features that embed long-range temporal dependencies. Dense LSTM[36], by contrast, is capable of capturing temporal dependencies in sequential data as memory cells can store information over a long period, allowing them to remember important information from earlier time steps. This capacity, nonetheless, is only partial owing to the shallow learning implemented by LSTM. Transformers, thanks to their self-attention mechanism for capturing dependencies between all positions in the input sequence, are much more effective in modeling long-term dependencies without the need for recurrent connections. On the other side, the self attention in frequency domain learns global dependencies for features representation. By leveraging conjointly long-range time-dependencies and global dependencies, our EmMixFormer is much more powerful to model the eye movement data w.r.t other approaches, which explains the new state-of-the-art results it achieves on the data collected at long or short intervals. Besides, Table IX shows that the performance of all approaches is degraded with increasing time duration. This is attributed to the fact that there are more variations in collected data with long time intervals, which results in increasing mismatching during recognition.
III-F Ablation experiment
Our our EmMixFormer consists of siamese CNN and Mix block, where Mix block includes three modules: Attention LSTM, Transformer, and Fourier Transformer (as shown in Fig.2). To investigate the effect of each module w.r.t recognition accuracy improvement, we conducted ablation experiments on the round 1 data from the RAN sub-dataset in GazeBase. To ease the description, we remove Mix block from EmMixFormer to obtain the basic siamese CNN. Then, we gradually add attention LSTM, Transformer, and Fourier Transformer, with the resulting models denoted as Siamese CNN+Transformer, Siamese CNN+LSTM+Transformer and Siamese CNN+attLSTM+Transformer, respectively. Afterwards, the Fourier Transformer, is included to the attLSTM+Transformer, with the resulting model represented as Siamese CNN+attLSTM+Transformer+Fourierformer, i.e. EmMixFormer. The verification error rates of these ablation schemes are shown in Table X and Fig.8. The experimental results demonstrate that incorporating Transformer into siamese CNN increases the accuracy, which implies that Transformer can learn dependency features to improve recognition performance. Also, we can observe that the combination of the Attention LSTM and the Fourier Transformer allows to improve the identification accuracy of simple Transformer. Besides, comparing Transformer with LSTM+Transformer, we observe that the LSTM can provide compensation features to improve the performance of Transformer, which is supported by the work [56]. From Table X and Fig.8, we also observe that attLSTM+Transformer achieves lower EER than LSTM+Transformer, which implies that the attention mechanism improves LSTM performance. Finally, the performance becomes optimal when attLSTM+Transformer is combined with Fourier Transformer, which demonstrates that the attention mechanism in the frequency domain captures global dependencies for feature representation.
TABLE X: Results of the ablation experiments on the RAN database
Method EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
Siamese CNN 0.1594 0.2498 0.7091 0.9900
+Transformer 0.1214 0.1491 0.6167 0.9133
LSTM+Transformer 0.0880 0.0776 0.4148 0.8204
attLSTM+Transformer 0.0936 0.0887 0.3720 0.7377
EmMixformer 0.0801 0.0680 0.2818 0.6087
Figure 8: ROC Curve of ablation experiment on RAN in the GazeBase dataset
IV Conclusion
This paper proposes an end-to-end model, termed EmMixFormer, for eye movement biometric authentication. EmMixFormer combines Attention LSTM, Transformer, and Fourier transformer for feature representation. As the attention LSTM and Transformer are capable of learning temporal dependencies within eye movement data while Fourier transformer can capture global feature dependencies, our proposed EmMixFormer is able to infer feature representation embeddings that combine long-range time-dependencies and global dependencies. Our experimental results on three representative datasets show that our approach outperforms the existing approaches and achieves a new state-of-the-art verification accuracy.
In the future, we will focus on the following research directions. First, we will collect more data to build larger datasets for evaluating eye-tracking models. This will provide a broader and more diverse dataset to improve performance. Second, we will refine the preprocessing methods and model architectures for eye-tracking data to enhance feature representation. Additionally, since eye movements are independent of existing biometric modalities, we can explore integrating eye movements as a supplementary biometric feature with other modalities. This fusion scheme can be leveraged in multimodal biometric systems to enhance system recognition accuracy and security.